# A functional limit theorem for a dynamical system with an observable maximised on a Cantor set

### v1

@@ 1.1. Enriched functional limit theorems
where $10<\theta<$ 1 is the extremal index, which gives the reciprocal of the average number of abnormal observations within a cluster and, in this case, is related to the decay rate of the measure of the iterative approximations of the maximal (Cantor) set, during its construction. In the case of the ternary Cantor set, $\theta=2/3$.
@@ 1.3. Structure of the paper
In Section 2 we outline our class of interval maps and the observables we will use, which are maximised on a dynamically defined Cantor set. In Section 3 we explain the theory of functional limit theorems for heavy tailed systems with clustering. This involved defining sequence spaces $l_{\infty}^{+},\tilde{l}_{\infty}^{+}$ aligned to a ‘block structure’, defining the anchored tail process, defining the dependence conditions Д$\D_{q_{n}}^{*}$ and Д$\D_{q_{n}}^{{}^{\prime}*}$, defining Rare Events Point Processes and explaining what the enriched functional limit theorem in [FFT25] is. In Section 4 we apply this theory to our class of examples. This long section is broken down into the statement of the main theorems, defining our constants and normalisation sequences, the extremal index, checking the dependence requirements, showing the anchored tail process is well defined and finally dealing with the small jumps condition in the case $21<\alpha<$ 2, using ideas from [CNT25].
@@ 2. A Fractal Maximal Set
Let $\Sigma=\Sigma_{\hat{k}}=\{1,\ldots,\hat{k}\}^{\mathbb{N}}$ and let $\sigma:\Sigma\to\Sigma$ be the left-shift map. For $x\in[0,1]$, let $\underline{x}=(x_{1},x_{2}\ldots)\in\Sigma$ be defined by $x_{i}=k\in\{1,\ldots,\hat{k}\}$ if $G^{i-1}(x)\in(I,J)_{k}$. Observe that $x\in\Lambda\iff\underline{x}=(a_{1},a_{2},\ldots)$ where $a_{i}\in\mathcal{I}_{I}$ for all $i\in\mathbb{N}$.
@@
Let $\underline{x}=(a_{1},a_{2},\ldots)\in\Sigma$. Let $\mathcal{I}_{x}=\{i\in\mathbb{N}:a_{i}\in\mathcal{I}_{J}\}$. Let
@@
$\psi$ provides us with a notion of distance to $\Lambda$ such that the earlier a point exits the dynamical construction of $\Lambda$ (in the sense of (2.1)) the bigger the corresponding distance to $\Lambda$.
@@
    $\{p_{1},\dots,p_{k_{I}^{n}}\}=\left\{x\in[0,1]:x_{i}\in\mathcal{I}_{I}\text{ for }i=1,\ldots n,\text{ and }x_{i}\in\mathcal{I}_{J}\text{ for }i\geq n+1\right\};$
@@ 3. General theory of functional limit theorems
We define sequences $(k_{n})_{n\in\mathbb{N}}$, $(r_{n})_{n\in\mathbb{N}}$, $(t_{n})_{n\in\mathbb{N}}$, which we assume to be such that
$k_{n},r_{n},t_{n}\xrightarrow[n\to\infty]{}\infty\quad\mbox{and}\quad k_{n}\cdot t_{n}=o(n),k_{n}\cdot r_{n}\sim n.$ (3.1)
We assume for each $\tau\geq 0$ there is $(u_{n}(\tau))_{n}$ such that
@@ 3.1. Sequence spaces and the anchored tail process
$l_{\infty}^{+}$ :=$:=\left\{\mathbf{x}=(x_{j})_{j}\in(0,\infty]^{\mathbb{Z}}\colon\;\lim_{|j|\to\infty}x_{j}=\infty\right\}.$
@@
We let ~$\tilde{\infty}\in\tilde{l}_{\infty}^{+}$ be the sequence where all terms are $\infty$.
@@
Here $(r_{n})_{n}$ is assumed to satisfy (3.1): in our applications it is the sequence appearing in Д$\D_{q_{n}}^{*}$ and Д$\D^{{}^{\prime}*}_{q_{n}}$ below. We call this the transformed tail process.
@@
3.2. The Д$\D_{q_{n}}^{*}$ and Д$\D_{q_{n}}^{{}^{\prime}*}$ conditions and the extremal index
For $B\subset[0,1]$, $J$ a subset of $[0,\infty)$ and $q\in\mathbb{N}$ define
@@ 3.2. The Д$\D_{q_{n}}^{*}$ and Д$\D_{q_{n}}^{{}^{\prime}*}$ conditions and the extremal index
Now let $E$ be an element of the ring generated by the half open intervals of $[0,1]$ so that $E=\cup_{\ell=1}^{N}J_{\ell}=\cup_{\ell=1}^{N}[a_{\ell},b_{\ell})$ and for each $\ell=1,\ldots,N$, let $A_{\ell}$ be an element of the ring generated by the half open intervals of $[0,\infty)$ so that $A_{\ell}=\cup_{s=1}^{N_{\ell}}[{\tau^{\prime\prime}}_{s},\tau^{\prime}_{s})$.
@@
Condition Д$\D_{q_{n}}^{*}$. We say that Д$\D_{q_{n}}^{*}$ holds for the sequence ${X}_{0},{X}_{1},\dots$ if there exist sequences $(k_{n})_{n\in\mathbb{N}},(r_{n})_{n\in\mathbb{N}},(t_{n})_{n\in\mathbb{N}}$ and $(q_{n})_{n\in\mathbb{N}}$ as defined in (3.1), such that for every $N,t,n\in\mathbb{N}$ and every $J_{\ell},A_{\ell}$ as above, for $\ell=1,\dots,N$, we have
$\left\lvert\mathbb{P}\left(A_{n,\ell}^{(q_{n})}\cap\bigcap_{i=\ell}^{N}\mathscr{W}_{J_{n,i}}\left(A_{n,i}^{(q_{n})}\right)\right)-\mathbb{P}\left(A_{n,i}^{(q_{n})}\right)\mathbb{P}\left(\bigcap_{i=\ell}^{N}\mathscr{W}_{J_{n,i}}\left(A_{n,i}^{(q_{n})}\right)\right)\right\rvert\leq\gamma(n,t_{n})$
@@
Condition Д$\D_{q_{n}}^{{}^{\prime}*}$. We say that Д$\D_{q_{n}}^{{}^{\prime}*}$ holds for the sequence ${X}_{0},{X}_{1},\dots$ if there exist sequences $(k_{n})_{n\in\mathbb{N}},(r_{n})_{n\in\mathbb{N}},(t_{n})_{n\in\mathbb{N}}$ and $(q_{n})_{n\in\mathbb{N}}$ as defined in (3.1), such that for every $A_{1}=\cup_{s=1}^{N_{1}}[{\tau^{\prime\prime}}_{s},\tau^{\prime}_{s})$,
@@ 3.3. Complete convergence of the Rare Events Point Processes
The complete convergence of these REPP follows from [FFT25, Theorem 3.18], which assuming that the transformed anchored tail process exists and is well defined (see Definition 3.1) and that conditions Д$\D_{q_{n}}^{*}$ and Д$\D_{q_{n}}^{*^{\prime}}$ are satisfied, states that $N_{n}$ given in (3.7) converges weakly in the space of boundedly finite point measures on $\mathbb{R}_{0}^{+}\times\tilde{l}_{\infty}^{+}\setminus\{\tilde{\infty}\}$ with weak# topology (see [FFT25, Appendix C] for details) to the Poisson process with decorations, which can be written as:
@@
where $T_{i}$ and $U_{i}$ are independently (of $i$ and of each other) distributed according to $Leb$ and $\theta Leb$ respectively, for $\theta$ as in (3.6), and $(\mathbf{\tilde{Q}}_{i})_{i}$ is an i.i.d sequence where for each $i$, $(\mathbf{\tilde{Q}}_{i,j})_{j}$ has the same distribution as $(Q_{j})_{j}$ in (3.4).
@@ 3.4. Enriched functional limit theorem
where the sequence $(c_{n})_{n\in\mathbb{N}}$ is such that $c_{n}=0$ if $10<\alpha<$ 1 and
$c_{n}=\frac{n}{\mathfrak{a}_{n}}\mathbb{E}\left(X_{0}{\mathbbm{1}}_{|X_{0}|\leq a_{n}}\right),\qquad\text{for }1<\alpha<2.$ (3.9)
@@
From [FFT25, Theorem 2.5], in the case $10<\alpha<$ 1, we obtain that the càdàg function $S_{n}(t)$ embedded in the space $F^{\prime}$ converges to $(V,\text{disc}(V),(e_{V}^{s})_{s\in\text{disc}(V)})$, where $V$ is an $\alpha$-stable Lévy process on $[0,1]$, which can be written as:
@@
In the case $21<\alpha<$ 2, we need the next two extra conditions to hold. Namely, for all $\delta>0$
@@ 4.1. Statement of main results
$N_{n}=\sum_{i=1}^{\infty}\delta_{\left(\frac{i}{k_{n}},\tilde{\pi}\left(\left(n(X_{(i-1)r_{n}})^{-\alpha},n(X_{(i-1)r_{n}+1})^{-\alpha},\ldots,n(X_{r_{n}-1})^{-\alpha}\right)\right)\right)},$
@@
where $c_{n}=0$, when $10<\alpha<$ 1 and is given by (3.9), when $21<\alpha<$ 2.
Then, the process $S_{n}(t)$ converges in $F^{\prime}$ to $(V,\text{disc}(V),(e_{V}^{s})_{s\in\text{disc}(V)})$, where $V$ is as described in (3.10), when $10<\alpha<$ 1, or as in (3.14), when $21<\alpha<$ 2, and the excursions can be written as
@@ 4.2. Tuning constants and normalising sequences
We consider that $j_{n,\tau}^{\ell}$ is infinite, whenever $a_{i}=0$ for all $i\geq\ell$ and in that case the same applies to all $j_{n,\tau}^{i}$, with $i\geq\ell$. Note that Lemma 2.2(c)(ii) implies that $j_{n,\tau}$ grows logarithmically in $n$. Moreover, by Lemma 2.4,
@@ 4.3. Extremal Index
We define
$\begin{split}U_{n}^{(q_{n})}(\tau)&:=\{X_{0}(x)>u_{n}(\tau),X_{1}\leq u_{n}(\tau),\dots,X_{q_{n}}\leq u_{n}(\tau)\}\\ &=\{\psi<u_{n,\tau},\psi\circ G\geq u_{n,\tau},\dots,\psi\circ G^{q_{n}}\geq u_{n,\tau}\}\end{split}$
which, for $n$ sufficiently large, can be rewritten
$\begin{split}U_{n}^{(q_{n})}(\tau)&=\{\psi<u_{n,\tau},\psi\circ G\geq u_{n,\tau}\}\\ &=\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})}\cap G^{-1}([0,1]\setminus(\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})}))\\ &=H^{(j_{n,\tau})}\cup(\Lambda_{j_{n,\tau}}\cap G^{-1}((\Lambda_{j_{n,\tau}-1}\setminus\Lambda_{j_{n,\tau}})\setminus H^{(j_{n,\tau})}))\end{split}$ (4.4)
so, in particular, $q_{n}=1$ for all $n\in\mathbb{N}$.

### v2

@@ 1.1. Enriched functional limit theorems
where $0<\theta<1$ is the extremal index, which gives the reciprocal of the average number of abnormal observations within a cluster and, in this case, is related to the decay rate of the measure of the iterative approximations of the maximal (Cantor) set, during its construction. In the case of the ternary Cantor set, $\theta=2/3$.
@@ 1.3. Structure of the paper
In Section 2 we outline our class of interval maps and the observables we will use, which are maximised on a dynamically defined Cantor set. In Section 3 we explain the theory of functional limit theorems for heavy tailed systems with clustering. This involved defining sequence spaces $l_{\infty}^{+},\tilde{l}_{\infty}^{+}$ aligned to a ‘block structure’, defining the anchored tail process, defining the dependence conditions $\D_{q_{n}}^{*}$ and $\D_{q_{n}}^{{}^{\prime}*}$, defining Rare Events Point Processes and explaining what the enriched functional limit theorem in [FFT25] is. In Section 4 we apply this theory to our class of examples. This long section is broken down into the statement of the main theorems, defining our constants and normalisation sequences, the extremal index, checking the dependence requirements, showing the anchored tail process is well defined and finally dealing with the small jumps condition in the case $1<\alpha<2$, using ideas from [CNT25].
@@ 2. A Fractal Maximal Set
Let $\Sigma=\Sigma_{\hat{k}}=\{1,\ldots,\hat{k}\}^{\mathbb{N}}$ and let $\sigma:\Sigma\to\Sigma$ be the left-shift map. For $x\in[0,1]$, let $\underline{x}=(x_{1},x_{2}\ldots)\in\Sigma$ be defined by $x_{i}=k\in\{1,\ldots,\hat{k}\}$ if $G^{i-1}(x)\in(I,J)_{k}$. Observe that $x\in\Lambda\iff\underline{x}=(x_{1},x_{2},\ldots)$ where $x_{i}\in\mathcal{I}_{I}$ for all $i\in\mathbb{N}$.
@@
Let $\underline{x}=(x_{1},x_{2},\ldots)\in\Sigma$. Let $\mathcal{I}_{x}=\{i\in\mathbb{N}:x_{i}\in\mathcal{I}_{J}\}$. Let
@@
$\psi$ provides us with a notion of distance to $\Lambda$ such that the earlier a point exits the dynamical construction of $\Lambda$ (in the sense of (2.1)) the bigger the corresponding distance to $\Lambda$. We will also view $\psi$ is a $[0,1]$-valued random variable defined on $([0,1],\mathcal{B}[0,1],\mathfrak{m})$.
@@
    $\{p_{1},\dots,p_{k_{I}^{n}}\}=\left\{x\in[0,1]:x_{i}\in\mathcal{I}_{I}\text{ for all }i=1,\ldots n,\text{ and }x_{i}\in\mathcal{I}_{J}\text{ for }i\geq n+1\right\};$
@@ 3. General theory of functional limit theorems
We define sequences $(k_{n})_{n\in\mathbb{N}}$, $(r_{n})_{n\in\mathbb{N}}$, $(t_{n})_{n\in\mathbb{N}}$, $(q_{n})_{n}$, which we assume to be such that
$k_{n},r_{n},t_{n}\xrightarrow[n\to\infty]{}\infty\quad\text{ where }k_{n}\cdot t_{n}=o(n),k_{n}\cdot r_{n}\sim n\text{ and }q_{n}=o(r_{n}).$ (3.1)
As in [FFT25, (3.2), (3.3)], these relate to a blocking argument with $k_{n}$ blocks of size $\lfloor n/k_{n}\rfloor$, where $t_{n}$ is a time gap inserted to achieve asymptotic independence of the blocks, while $q_{n}$ is the clustering run length, i.e., all abnormal observations occurring within a time difference of at most $q_{n}$ units between each other belong to the same cluster. In most dynamical applications $q_{n}=q$ for all $n\in\mathbb{N}$, such as when the observable is maximised at a single periodic point of period $p$, in which case $q_{n}=p$ (see [FFT12]). In fact, we will see that, in this paper, for our applications we only require $q_{n}=1$. We assume for each $\tau\geq 0$ there is $(u_{n}(\tau))_{n}$ such that
@@ 3.1. Sequence spaces and the anchored tail process
$l_{\infty}^{+}$ $:=\left\{\mathbf{x}=(x_{j})_{j}\in(0,\infty]^{\mathbb{Z}}\colon\;\lim_{|j|\to\infty}x_{j}=\infty\right\}.$
@@
We let $\tilde{\infty}\in\tilde{l}_{\infty}^{+}$ be the sequence where all terms are $\infty$.
@@
Here $(r_{n})_{n}$ is assumed to satisfy (3.1): in our applications it is the sequence appearing in $\D_{q_{n}}^{*}$ and $\D^{{}^{\prime}*}_{q_{n}}$ below. We call this the transformed tail process.
@@
3.2. The $\D_{q_{n}}^{*}$ and $\D_{q_{n}}^{{}^{\prime}*}$ conditions and the extremal index
For $B\subset[0,1]$, $J$ a subset of $0,\infty)$ and $q\in\mathbb{N}$ define
@@ 3.2. The $\D_{q_{n}}^{*}$ and $\D_{q_{n}}^{{}^{\prime}*}$ conditions and the extremal index
Now let $E$ be an element of the ring generated by the half open intervals of $[0,1]$ so that $E=\cup_{\ell=1}^{N}J_{\ell}=\cup_{\ell=1}^{N}[a_{\ell},b_{\ell})$ and for each $\ell=1,\ldots,N$, let $A_{\ell}$ be an element of the ring generated by the half open intervals of $0,\infty)$ so that $A_{\ell}=\cup_{s=1}^{n_{\ell}}[{\tau_{s}^{\prime\prime}},\tau^{\prime}_{s})$ for some $n_{\ell}\in\mathbb{N}$ and some $\tau_{1}^{\prime\prime}<\tau_{1}^{\prime}<\tau_{2}^{\prime\prime}<\cdots<\tau_{n_{\ell}}^{\prime\prime}<\tau_{n_{\ell}}^{\prime}$.
@@
Condition $\D_{q_{n}}^{*}$. We say that $\D_{q_{n}}^{*}$ holds for the sequence ${X}_{0},{X}_{1},\dots$ if there exist sequences $(k_{n})_{n\in\mathbb{N}},(r_{n})_{n\in\mathbb{N}},(t_{n})_{n\in\mathbb{N}}$ and $(q_{n})_{n\in\mathbb{N}}$ as defined in (3.1), such that for every $N,t,n\in\mathbb{N}$ and every $J_{\ell},A_{\ell}$ as above, for $\ell=1,\dots,N$, we have
$\left\lvert\mathbb{P}\left(A_{n,\ell}^{(q_{n})}\cap\bigcap_{i=\ell}^{N}\mathscr{W}_{J_{n,i}}\left(A_{n,i}^{(q_{n})}\right)\right)-\mathbb{P}\left(A_{n,\ell}^{(q_{n})}\right)\mathbb{P}\left(\bigcap_{i=\ell}^{N}\mathscr{W}_{J_{n,i}}\left(A_{n,i}^{(q_{n})}\right)\right)\right\rvert\leq\gamma(n,t_{n})$
@@
Condition $\D_{q_{n}}^{{}^{\prime}*}$. We say that $\D_{q_{n}}^{{}^{\prime}*}$ holds for the sequence ${X}_{0},{X}_{1},\dots$ if there exist sequences $(k_{n})_{n\in\mathbb{N}},(r_{n})_{n\in\mathbb{N}},(t_{n})_{n\in\mathbb{N}}$ and $(q_{n})_{n\in\mathbb{N}}$ as defined in (3.1), such that for every $A_{1}=\cup_{s=1}^{N_{1}}[{\tau^{\prime\prime}}_{s},\tau^{\prime}_{s})$,
@@
Remark 3.2.
Note that if $(q_{n})_{n\in\mathbb{N}}$ and $(\tilde{q}_{n})_{n\in\mathbb{N}}$ are two sequences satisfying (3.1) and $q_{n}\leq\tilde{q}_{n}$, for all $n\in\mathbb{N}$. Then $\D^{\prime}_{q_{n}}$ implies $\D^{\prime}_{\tilde{q}_{n}}$. As observed in [AFF20, Remark 2.11], one should try first to find the smallest constant sequence $q_{n}=q$ that makes $\D_{q_{n}}^{{}^{\prime}*}$ hold, establishing, in this way, the run length. We will see that here we can actually take $q_{n}=1$ for all $n\in\mathbb{N}$.
@@ 3.3. Complete convergence of the Rare Events Point Processes
The complete convergence of these REPP follows from [FFT25, Theorem 3.18], which assuming that the transformed anchored tail process exists and is well defined (see Definition 3.1) and that conditions $\D_{q_{n}}^{*}$ and $\D_{q_{n}}^{*^{\prime}}$ are satisfied, states that $N_{n}$ given in (3.7) converges weakly in the space of boundedly finite point measures on $\mathbb{R}_{0}^{+}\times\tilde{l}_{\infty}^{+}\setminus\{\tilde{\infty}\}$ with weak# topology (see [FFT25, Appendix C] for details) to the Poisson process with decorations, which can be written as:
@@
where $T_{i}$ and $U_{i}$ are independently (of $i$ and of each other) distributed according to $Leb$ and $\theta Leb$ respectively, for $\theta$ as in (3.6), and $(\mathbf{\tilde{Q}}_{i})_{i\in\mathbb{N}}$ is an i.i.d sequence such that each $\mathbf{\tilde{Q}}_{i}\in\tilde{l}_{\infty}^{+}\setminus\{\tilde{\infty}\}$ has the same distribution as $\tilde{\pi}((Q_{j})_{j\in\mathbb{Z}})$, where the bi-infinite sequence $(Q_{j})_{j}$ is as in (3.4).
@@ 3.4. Enriched functional limit theorem
where the sequence $(c_{n})_{n\in\mathbb{N}}$ is such that $c_{n}=0$ if $0<\alpha<1$ and
$c_{n}=\frac{n}{\mathfrak{a}_{n}}\mathbb{E}\left(X_{0}{\mathbbm{1}}_{|X_{0}|\leq\mathfrak{a}_{n}}\right),\qquad\text{for }1<\alpha<2.$ (3.9)
@@
From [FFT25, Theorem 2.5], in the case $0<\alpha<1$, we obtain that the càdàg function $S_{n}(t)$ embedded in the space $F^{\prime}$ converges to $(V,\text{disc}(V),(e_{V}^{s})_{s\in\text{disc}(V)})$, where $V$ is an $\alpha$-stable Lévy process on $[0,1]$, which can be written as:
@@
In the case $1<\alpha<2$, we need the next two extra conditions to hold. Namely, for all $\delta>0$
@@ 4.1. Statement of main results
$N_{n}=\sum_{i=1}^{\infty}\delta_{\left(\frac{i}{k_{n}},\tilde{\pi}\left(\left(n(X_{(i-1)r_{n}})^{-\alpha},n(X_{(i-1)r_{n}+1})^{-\alpha},\ldots,n(X_{ir_{n}-1})^{-\alpha}\right)\right)\right)},$
@@
where $c_{n}=0$, when $0<\alpha<1$ and is given by (3.9), when $1<\alpha<2$.
Then, the process $S_{n}(t)$ converges in $F^{\prime}$ to $(V,\text{disc}(V),(e_{V}^{s})_{s\in\text{disc}(V)})$, where $V$ is as described in (3.10), when $0<\alpha<1$, or as in (3.14), when $1<\alpha<2$, and the excursions can be written as
@@ 4.2. Tuning constants and normalising sequences
If, for some $\ell>1$, we have $a_{i}=0$ for all $i\geq j_{n,r}^{\ell}$, then we set $j_{n,r}^{s}=\infty$ for all $s\geq\ell$. Note that Lemma 2.2(c)(ii) implies that $j_{n,\tau}$ grows logarithmically in $n$ (i.e., , $\limsup_{n\to\infty}\frac{j_{n,\tau}}{\log n}<\infty$). Moreover, by Lemma 2.4,
@@ 4.3. Extremal Index
Recall that
$\begin{split}U_{n}^{(q)}(\tau)&=\{X_{0}(x)>u_{n}(\tau),X_{1}\leq u_{n}(\tau),\dots,X_{q}\leq u_{n}(\tau)\}\\ &=\{\psi<u_{n,\tau},\psi\circ G\geq u_{n,\tau},\dots,\psi\circ G^{q}\geq u_{n,\tau}\}.\end{split}$
Observe that for $n$ sufficiently large, we have
$\begin{split}U_{n}^{(1)}(\tau)&=\{\psi<u_{n,\tau},\psi\circ G\geq u_{n,\tau}\}\\ &=\left(\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})}\right)\cap G^{-1}([0,1]\setminus(\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})}))\\ &=H^{(j_{n,\tau})}\cup(\Lambda_{j_{n,\tau}}\cap G^{-1}((\Lambda_{j_{n,\tau}-1}\setminus\Lambda_{j_{n,\tau}})\setminus H^{(j_{n,\tau})}))\end{split}$ (4.4)
Now, observing that the points of $U_{n}^{(1)}(\tau)$ are sent by $G$ into the holes of smaller and smaller generation in the construction of $\Lambda$, it follows that $U_{n}^{(1)}(\tau)=U_{n}^{(2)}(\tau)=U_{n}^{(3)}(\tau)=\ldots$. In light of Remark 3.2, this means that $q_{n}=1$ is then a natural choice to prove condition $\D_{q_{n}}^{{}^{\prime}*}$, which is indeed established in Section 4.4.